# Distilling Machine Learning’s Added Value: Pareto Fronts in Atmospheric Applications

### v1

@@ 4.2.4 Distilling Radiative Transfer’s Bidirectionality
Following Ukkonen (2022), we divide the shortwave radiative fluxes by the incoming TOA downward flux (), to implicitly inform the NN that the Sun is the original energy source. At each vertical level , we then target the resulting ratios ( and ) with a bidirectional RNN. The RNN has two 64-unit gated recurrent unit layers (one forward, one backward), which are concatenated before a dense layer for predicting the fluxes, resulting in only 54,786 trainable parameters. The RNN includes a multiplication layer with and a custom layer to directly calculate shortwave heating rates, maintaining the same optimization objective as the other models. We find that the resulting RNN (green star) is clearly Pareto-optimal, competing with the U-net++ and achieving MSE below 0.1 K2 day-2, even in the complex, multi-cloud case. Our results highlight the advantages of physically constraining ML solutions using robust knowledge of the underlying system, extending the aquaplanet findings of Bertoli et al. (2023) to an Earth-like, operational setting, and paving the way towards fully hybrid physics-ML emulators of shortwave radiation (Schneiderman, 2024). We now turn to a problem in which both spatial and temporal connectivities may be considered: precipitation parameterization.
@@ 4.3.1 Motivation
where represents a continuous, bidimensional horizontal coordinate. This spatial averaging removes convective organization information below the coarse grid’s spatial scale. Using storm-resolving simulations as a reference, Shamekh et al. (2023) showed this blurs precipitation, adding irreducible uncertainty, especially for large values of precipitable water (PW). Given that high-resolution information is typically inaccessible in a parameterization context, we ask in this section: How much of this lost spatial granularity can we recover with temporal memory?
@@ 4.3.2 Setup
To address this question, we build upon the framework of Shamekh et al. (2023), coarse-graining SAM-DYAMOND (Khairoutdinov et al., 2022) simulations from their native, high-resolution horizontal grid (4.34-km spacing at the equator) to a low-resolution grid (138.9-km spacing at the equator) representative of a coarse Earth system model. Over the tropical ocean (S-N), we map 6 environmental variables – PW, 2-m specific humidity [kg kg-1], 2-m temperature [K], surface sensible and latent heat fluxes [], and sea surface temperature [K] – to 15-min precipitation rates [mm h-1]. We use three setups at a given low-resolution location including the high-resolution grid locations :
where PW′ is PW’s anomaly with respect to its grid cell-mean and memory is accounted for through up to two time steps in the past, in which case . Samples are extracted from 10 days of the simulation after spin-up, with 6 days for training, 2 for validation, and 2 for testing. To focus on precipitation intensity rather than triggering, samples with precipitation values below 0.05 mm/h are discarded, resulting in a total of samples.
4.3.3 Comparing Spatial and Temporal Connectivities
As illustrated in Fig. 4, we design three categories of NNs to learn the three mappings given by Eq. 28, 29, and 30. First, our baseline NNs (orange crosses) are simple MLPs with 5 layers of different widths, resulting in a number of trainable parameters ranging from and never achieving MSEs below 0.4 mm2 h-2. The NNs using spatial granularity (purple crosses) additionally encode the high-resolution PW anomaly field through an encoder-decoder structure with a small bottleneck representing convective organization as a bidimensional latent variable, optionally regularized via data augmentation applied to PW′. MSEs are below 0.04 mm2 h-2 for the deepest encoder-decoder architectures, which yield high reconstruction quality through more than 1M trainable parameters. The NNs leveraging temporal memory instead of spatial resolution (green crosses) use the last or last two previous timesteps to achieve competitive MSEs, all below 0.15 mm2 h-2. These “Memory-NNs” are built using two types of architectures: On the one hand, we flatten previous timesteps and feed them to MLPs conditioned on non-zero current precipitation, which ignores temporal connectivity. On the other hand, we use CNN-based temporal models that preserve temporal ordering through 1D convolutional layers, achieving MSEs around 0.1 mm2 h-2 with times fewer trainable parameters than the “High-Res Inputs-NNs” that use storm-scale information. Overall, our results showcase the added value of spatial and temporal connectivities, as while:
@@ 4.3.4 Mitigating Low Spatial Resolution with Memory
First, we can ask: Why would temporal memory of the coarse environment be relevant on its own, given that at a given time , the precipitation can be exactly diagnosed from the high-resolution environment ? The informativeness of high spatial resolution is confirmed by the fact that our lowest-error model is . As written in Eq. 29, decomposes the high-resolution information into a low-resolution component and the inputs’ spatial anomaly . Using this decomposition, we can write the following causal graph:
@@
where we posit that for a small enough time step , the combination of the high-resolution anomaly and low-resolution information is enough to diagnose precipitation and step the entire system forward in time from to . Formally, this is a Markovian assumption and eliminates the need for temporal memory. Using Eq. 32 at times and , we see that in this framework, suppressing access to high-resolution information “unblocks” the following causal path:
@@
creating a conditional dependence of precipitation on the past low-resolution environment by removing their d-separation given current high-resolution anomalies (Pearl et al., 2000). This explains the need for temporal memory without high-resolution information and provides a rationale for the Pareto optimality of the “simple-memory NN” that considers the low-resolution inputs and their past timesteps.
@@
Now, we can ask: How much information is lost in the process of replacing the high-resolution spatial anomaly with temporal memory of the low-resolution spatial means? The answer depends on the distance between low and high resolutions. There is no difference if these resolutions are equal, while we do not expect low-resolution memory to help at all if the low-resolution grid is coarse to the point that we cannot distinguish very different high-resolution situations. For the resolutions considered here, the bidimensionality of the encoder-decoder’s bottleneck suggests that the high-resolution anomaly field can be represented by a bidimensional variable for the purpose of modeling precipitation: , which represents the aspects of convective organization that explain why precipitation can be different (stochastic) for the same low-resolution inputs (Moseley et al., 2016). This simplifies the question to how well we can represent from the time series of low-resolution inputs. The encouraging results of our models leveraging temporal memory, along with the finding of Shamekh et al. (2023) that is accurately predicted by an autoregressive model informed by the coarse-scale environment’s past timesteps, suggest a mostly positive response. This reassuringly confirms that Earth system models with limited spatial resolutions can realistically represent coarse-scale precipitation as long as efforts to improve precipitation parameterization continue (Schneider et al., 2024). More broadly, this confirms that Pareto-optimal model hierarchies are useful in empirically establishing the partial equivalence between temporal memory and spatial granularity. This has practical applications for multi-scale dynamical systems that are not self-similar, where ergodicity cannot be used to deterministically infer detailed spatial information from coarse time series data.
@@ 5 Conclusion
In this study, we demonstrated that Pareto-optimal model hierarchies within a well defined (complexity, error) plane not only guide the development of data-driven models—from simple equations to sophisticated neural networks—but also promote process understanding. To distill knowledge from these hierarchies, we propose a multi-step approach: (1) use models along the Pareto front to hierarchically investigate the added value (functional representation, feature assimilation, spatial and temporal connectivities) that leads to incremental error decreases from the simplest to the most complex Pareto-optimal model; (2) generate hypotheses and propose models tailored to the system of interest based on this added value; and (3) once the models are sparse enough to be interpretable, reveal the system’s underlying properties that previous theories or models may have overlooked. Beyond knowledge discovery, such hierarchies promote interpretability by explaining the added value of models that may initially seem overly complex, and sustainability by optimizing models’ computational costs once their added value is justified.

### v2

@@ 4.2.4 Distilling Radiative Transfer’s Bidirectionality
Following Ukkonen (2022), we divide the shortwave radiative fluxes by the incoming TOA downward flux (), to implicitly inform the NN that the Sun is the original energy source. At each vertical level , we then target the resulting ratios ( and ) with a bidirectional RNN. The RNN has two 64-unit gated recurrent unit layers (one forward, one backward), which are concatenated before a dense layer for predicting the fluxes, resulting in only 54,786 trainable parameters. The RNN includes a multiplication layer with and a custom layer to directly calculate shortwave heating rates, maintaining the same optimization objective as the other models. We find that the resulting RNN (green star) is clearly Pareto-optimal, competing with the U-net++ and achieving MSE below 0.1 K2 day-2, even in the complex, multi-cloud case. Our results highlight the advantages of physically constraining ML solutions using robust knowledge of the underlying system, extending the aquaplanet findings of Bertoli et al. (2023) to an Earth-like, operational setting, and paving the way towards fully hybrid physics-ML emulators of shortwave radiation (Schneiderman, 2024).
@@ 4.3.1 Motivation
where represents a continuous, bidimensional horizontal coordinate. This spatial averaging removes convective organization information below the coarse grid’s spatial scale. Using storm-resolving simulations as a reference, Shamekh et al. (2023) showed this blurs precipitation, adding irreducible uncertainty, especially for large values of precipitable water (). Given that high-resolution information is typically inaccessible in a parameterization context, we ask in this case study: How much of this lost spatial granularity can we recover with temporal memory?
@@ 4.3.2 Setup
To address this question, we build upon the framework of Shamekh et al. (2023), coarse-graining SAM-DYAMOND (Khairoutdinov et al., 2022) simulations from their native, high-resolution horizontal grid (4.34-km spacing at the equator) to a low-resolution grid (138.9-km spacing at the equator) representative of a coarse Earth system model. Over the tropical ocean (S-N), we map 6 environmental variables – precipitable water [mm], 2-m specific humidity [kg kg-1], 2-m temperature [K], surface sensible and latent heat fluxes [], and sea surface temperature [K] – to 15-min precipitation rates [mm h-1]. We use three setups at a given low-resolution location including the high-resolution grid locations :
where is the precipitable water’s anomaly with respect to its grid cell-mean and memory is accounted for through up to two time steps in the past, in which case . Samples are extracted from 10 days of the simulation after spin-up, with 6 days for training, 2 for validation, and 2 for testing. To focus on precipitation intensity rather than triggering, samples with precipitation values below 0.05 mm/h are discarded, resulting in a total of samples.
4.3.3 Model Hierarchy
As illustrated in Fig. 4, we design three categories of NNs to learn the three mappings given by Eq. 29, 30, and 31. First, our baseline NNs (orange crosses) are simple MLPs with 5 layers of different widths, resulting in a number of trainable parameters ranging from and never achieving MSEs below 0.4 mm2 h-2. The NNs using spatial granularity (purple crosses) additionally encode the high-resolution precipitable water anomaly field through an encoder-decoder structure with a small bottleneck representing convective organization as a bidimensional latent variable, optionally regularized via data augmentation applied to . MSEs are below 0.04 mm2 h-2 for the deepest encoder-decoder architectures, which yield high reconstruction quality through more than 1M trainable parameters. The NNs leveraging temporal memory instead of spatial resolution (green crosses) use the last or last two previous timesteps to achieve competitive MSEs, all below 0.15 mm2 h-2. These “Memory-NNs” are built using two types of architectures: On the one hand, we flatten previous timesteps and feed them to MLPs conditioned on non-zero current precipitation, which ignores temporal connectivity. On the other hand, we use CNN-based temporal models that preserve temporal ordering through 1D convolutional layers, achieving MSEs around 0.1 mm2 h-2 with times fewer trainable parameters than the “High-Res Inputs-NNs” that use storm-scale information. Overall, our results showcase the added value of spatial and temporal connectivities, as while:
@@ 4.3.4 Mitigating Low Spatial Resolution with Memory
First, we can ask: Why would temporal memory of the coarse environment be relevant on its own, given that at a given time , the precipitation can be exactly diagnosed from the high-resolution environment ? The informativeness of high spatial resolution is confirmed by the fact that our lowest-error model is . As written in Eq. 30, decomposes the high-resolution information into a low-resolution component and the inputs’ spatial anomaly (denoted by primes). Using this decomposition, we can write the following causal graph:
@@
where we posit that for a small enough time step , the combination of the high-resolution anomaly and low-resolution information suffices to diagnose precipitation and step the entire system forward in time from to . Formally, this is a Markovian assumption, which eliminates the need for temporal memory. Under this assumption, the precipitation is independent from and previous timesteps when conditioned on the current environment . Using the bottom two lines of Eq. 34, we see that suppressing access to high-resolution information activates the following causal path from to by introducing a dependency:
@@
This induces a conditional dependence of precipitation on the past low-resolution environment by removing d-separation through the exclusion of current high-resolution anomalies from the conditioning set (Pearl et al., 2000). Therefore, temporal memory becomes necessary in the absence of high-resolution information, consistent with the general result of the Mori-Zwanzig decomposition for dynamical systems (Lucarini and Chekroun, 2023). This explains the Pareto optimality of the ”simple-memory NN” that considers low-resolution inputs and their past timesteps.
@@
Now, we can ask: How much information is lost in the process of replacing the high-resolution spatial anomaly with temporal memory of the low-resolution spatial means? The answer depends on the distance between low and high resolutions. There is no difference if these resolutions are equal, while we do not expect low-resolution memory to help at all if the low-resolution grid is coarse to the point that we cannot distinguish very different high-resolution situations. For the resolutions considered here, the encoder-decoder architecture (purple crosses) achieves high performance with a bottleneck whose dimension is only two. This suggests that the high-resolution anomaly field can be represented by a bidimensional (latent) variable, which we denote , for the purpose of modeling precipitation. encapsulates the aspects of convective organization that explain why precipitation can be different (stochastic) for the same low-resolution inputs (Moseley et al., 2016). This simplifies the question to how well we can represent from the time series of low-resolution inputs. The encouraging results of our models leveraging temporal memory, along with the finding of Shamekh et al. (2023) that is accurately predicted by an autoregressive model informed by the coarse-scale environment’s past timesteps, suggest a mostly positive response. This reassuringly confirms that Earth system models with limited spatial resolutions can realistically represent coarse-scale precipitation as long as efforts to improve precipitation parameterization continue (Schneider et al., 2024). In hybrid modeling applications, where the NN itself computes precipitation at previous time steps, error propagation may reduce the informativeness of temporal memory—a limitation worth exploring in future work. More broadly, our results confirm that Pareto-optimal model hierarchies are useful in empirically establishing the partial equivalence between temporal memory and spatial granularity. This has practical applications for multi-scale dynamical systems that are not self-similar, where ergodicity cannot be used to deterministically infer detailed spatial information from coarse time series data.
@@ 5 Conclusion
In this study, we demonstrated that Pareto-optimal model hierarchies within a well defined (complexity, error) plane not only guide the development of data-driven models—from simple equations to sophisticated neural networks—but also promote process understanding. While these hierarchies are derived empirically based on the models and data at hand, they serve as practical approximations for exploring trade-offs between error and complexity. To distill knowledge from these hierarchies, we propose a multi-step approach: (1) use models along the Pareto front to hierarchically investigate the added value that leads to incremental error decreases from the simplest to the most complex Pareto-optimal model; (2) generate hypotheses and propose models tailored to the system of interest based on this added value; and (3) once the models are sparse enough to be interpretable, reveal the system’s underlying properties that previous theories or models may have overlooked. Beyond knowledge discovery, such hierarchies promote interpretability by explaining the added value of models that may initially seem overly complex, and sustainability by optimizing models’ computational costs once their added value is justified.